Use two equations in two variables to solve the application.
At a store, ice cream cones cost $1.90 and sundaes cost $2.65. One day, the receipts for a total of 142 cones and sundaes were $312.55. How many cones were sold?

Answer by addingup(3677)

I already did one. I'm going to do this one, you should try to do the others because we already know math and you are just now learning
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
ice cream cones: i
sundaes: s
i + s = 142 therefore i = 142 - s
1.90i + 2.65s = 312.55
1.90(142-s) + 2.65s = 312.55
269.80 - 1.90s + 2.65s = 312.55
0.75s = 42.75
s = 57
A total of 57 sundaes were sold
And 142-57 = 85 cones